OpenStudy (anonymous):

What sine function represents an amplitude of 4, a period of pi over 2, no horizontal shift, and a vertical shift of −3?

OpenStudy (anonymous):

put a 4 out front for starters

OpenStudy (anonymous):

vertical shif of \(-3\) means put a \(-3\) out at the end

OpenStudy (anonymous):

so \[4\sin(bx)-3\] to find \(b\) solve \[\frac{2\pi}{b}=\frac{\pi}{2}\]